Consider the line containing the points C(2,9) and D(−1,4).
A: What is the slope of the line?
B: What is the equation of the line in point-slope form?
Select one answer choice for question A, and select all correct answer choices for question B.
B: y−9=−5/3(x−2)
B: y−9=5//3(x−2)
B: y−4=35(x+1)
A: −3/5
A: 5/3
B: y+4=3/5(x−1)
B: y−4=5/3(x+1)
A: 3/5
B: y−9=5/3(x+2)
A: −5/3
B: y−9=3/5(x−2)
B: y+9=−3/5(x+2)

Answer by stanbon(75887)

Consider the line containing the points C(2,9) and D(−1,4).
A: What is the slope of the line?
slope = (9-4)/(2--1) = 5/3
-------------------------------
B: What is the equation of the line in point-slope form?
y-4 = (5/3)(x+1)
-------------------------------
Cheers,
Stan H.
-------------
Select one answer choice for question A, and select all correct answer choices for question B.
B: y−9=−5/3(x−2)
B: y−9=5//3(x−2)
B: y−4=35(x+1)
A: −3/5
A: 5/3
B: y+4=3/5(x−1)
B: y−4=5/3(x+1)
A: 3/5
B: y−9=5/3(x+2)
A: −5/3
B: y−9=3/5(x−2)
B: y+9=−3/5(x+2)